9. Ato Bekele is producing various biscuits in his Company in three factories residing in different parts of Ethiopia. Its line of biscuits exhibits a highly seasonal demand pattern, with peaks during the month of September & October and declines through the months November & December. Given the following costs and monthly sales forecasts, determine whether;
(a) Level production, or (b) chase demand would more economically meet the demand for biscuits:
Months Sales forecast
September 40,000
October 25,000
November 60,000
December 75,000
Hiring costs-------------------------------------------------$ 100 per worker
Firing cost---------------------------------------------------$ 500 per worker
Inventory carrying cost------------------------------------$ 0.5 Cents per biscuits per biscuits
Regular production cost per biscuit----------------------$ 2
Production per employee-------------------------------- 1000 biscuits per month
Beginning work force-------------------------------------100 workers
1
Expert's answer
2020-06-05T10:11:59-0400

Total production in all four months =40000+25000+60000+75000=200000

Total average production = 50,000

For optimum level we know that


EOQ= "\sqrt{\frac{2DA}{h}}"


EOQ="\sqrt{\frac{2\times 50000\times2 }{0.5}}=" 633


So, Level production will be economical to meet the demand